Hello, my question is: Howard has $2.05 in quarters and dimes. He has 4 more quarters than dimes. Find the number he has of each coin.
Thank you.

Found 2 solutions by mananth, oberobic:
Answer by mananth(16949)

dimes ---- x
quarters x + 4
Convert dollars to cents
Value= 205

10 x + 25 x + 100= 205
35 x + 100 = 205
35 x = 205 -100
35 x = 105
/ 35
x= 3 dimes 30
7 quarters

Answer by oberobic(2304)

With coin problems you have to keep track of the number of coins and the value of the coins.
.
d = number of dimes
10d = value of the dimes in cents
q = number of quarters
25q = value of the quarters in cents
.
10d cents + 25q cents = 205 cents
.
He has 4 more quarters than dimes
q = d+4
.
Substitute
.
10d + 25(d+4) = 205
10d + 25d + 100 = 205
35d = 105
d = 3
.
q = d+4
q = 7
.
Check the values to see if this solution is the correct answer.
10(3) = 30 cents
25(7) = 175 cents
30+175 = 205 cents
Correct.
.
Answer: 3 dimes and 7 quarters.
.
Done.